# Electronic structure and transport in materials with flat bands: 2D materials and quasicrystals

### v1

@@ 2.3.5 Heterostrain and atomic relaxation
The angle of rotation $\theta$ between the two graphene layers is the main parameter for the existence of flat moiré bands, however for angles close to the magic angle, other parameters are also very important and can strongly modify the band structure. Many studies have shown the importance of atomic relaxation [107, 108, 109, 110, 111, 112, 114, 115, 116, 117, 118, 119, 120, 121, 122, 247]. In particular, the modulation of the position of the atoms along the axis perpendicular to the graphene planes ($z$-modulation) [116] among other things modifies the presence of gaps isolating the four flat bands from the other bands. It has also been shown that atomic relaxation does not strongly modify the bands at the first magic angle, $\theta\approx 1^{\circ}$ (Sec. 2.3.3), while it has a much stronger effect at smaller angles, which can destroy the notion itself of magic angles beyond the first [119].
@@
However, in the case of TBG on SiC close to the first magic angle, a detailed study [104, 106] shows that the effect of “heterostrain,” i.e., an in-plane deformation of one layer with respect to the other, is more important in many samples than the atomic relaxation. It is now well established that the heterostrain is present in many samples of TBG, although not well controlled experimentally, and it has strong consequences for the electronic properties of flat bands [104, 105, 106, 215, 248, 249, 250, 251, 227, 228, 226, 231, 232]. To study it in a quantitative way, it is then possible to calculate the local electronic structure from our SK-TB model (Sec. 2.3.2), without other adjustable parameters. This approach has allowed to study several samples that have have been studied from STM images in the literature (see Refs. [104, 106, 252, 253] and Refs. therein). The main result is that for TBGs with angles close to the first magic angle, a small amount of heterostrain is an ubiquitous factor in strongly modifying the electronic structure of the low-energy flat bands. For example, the uniaxial heterostrain can induce the appearance of a third peak in the low-energy DOS in addition to two van Hove singularities [104]. Moreover heterostrain modifies also the electron-electron interactions in many samples that have Fermi levels in the flat bands [106]. It therefore appears that for the first magic-angle TBGs studied experimentally, the heterostrain is a determining parameter, more important than the atomic relaxations.
For angles smaller than the first magic angle, several studies show the importance of atomic reconstruction [119] with a reduction in AA zones in favor of AB zones, which are energetically more stable. For very large moiré of several millions of atoms, this atomic reconstruction is stronger and can lead to new phenomena such as giant swirls associated with new types of electronic localization [243].
@@ 2.3.6 Magnetism
Figure 10: MFT results for TBG: (a) Magnetization ($\mu_{B}$) at zero temperature ($T=0$) as a function of $U$ at rotation angles $\theta=1.08^{\circ}$ (bilayer (30,31), $N=11164$), $\theta=1.30^{\circ}$ ((25,26), $N=7804$), and $\theta=1.47^{\circ}$ ((22,23), $N=6076$), respectively. $N$ is the number of orbitals per moiré cell. The monolayer case is calculated with similar conditions (similar supercell). For a monolayer, the critial $U$ above which the solution is magnetic is $U_{Cm}\approx 2.09t_{0}$, where $t_{0}=2.7$ eV is the first intralayer coupling. The other panels are for $\theta=1.08^{\circ}$ TBG: (b) Magnetization per moiré cell: (solid line) $M_{tot}^{*}(U)$ for the AF and FERRO solutions, and (dashed line) $M_{tot}(U)$ for FERRO ($M_{tot}=0$ for AF), see Eq. (31). (c) $M^{*}_{tot}(U=t_{0}=2.7\,{\rm eV})$ versus temperature $T$ for the AF solution. (d) Up and down local DOSs on an orbital with up magnetization at the center of the AA stacking region for $U=t_{0}=2.7$ eV for the AF solution. (e) and (f) Bands around $E_{F}$ of the AF and FERRO solutions for $U=t_{0}=2.7$ eV. Up and down bands are degenerate for the AF solution. SK-TB model is the one which results in an intraband monolayer velocity $V_{MLG}=0.79\cdot 10^{6}$ ms-1 (Sec. 2.3.2). TBGs are rigid, i.e., without atomic relaxation. For details of the MFT calculations see Ref. [254].
The large reduction in kinetic energy in flat bands increases the relative importance of interactions, making bilayer systems much more susceptible to correlation effects than a single layer [255, 102, 256, 216, 217, 257, 258, 259, 260, 261, 262, 130, 263, 22] (see also [254] and Refs. therein). One example is the magnetization of edge states in graphene nanoribbons [264, 265, 266, 267, 268, 269, 270, 271, 272, 273, 274], TMD nanoribbons [275], or graphene nanoflakes [276, 277, 278, 268, 269, 273, 279, 280, 281, 282, 283]. However, in TBGs the flat bands at the origin of strong electronic correlations are not due to defects or edges, but to the particular symmetry of the moiré pattern. We have first investigated [254] the question of magnetic instabilities at half filling of TBG using our tight-binding description of the non-interacting bilayer systems (Eq. (15)) to which we add an on-site Hubbard interaction $U$ in order to model the Coulomb repulsion between electrons of opposite spin [284],
@@
where $\sigma=\,\uparrow$ ($\downarrow$) is the spin up (down), $\hat{H}_{0\sigma}$ is the non-interacting SK-TB Hamiltonian for each spin (Eq. (19)), and $n_{i\sigma}=c_{i\sigma}^{{\dagger}}c_{i\sigma}$ is the number operator at site $i$ and $\mu_{c}$ the chemical potential. In the static mean-field (MFT) approximation, assuming that local magnetic moments are collinear to the $z$-direction (Hartree approximation), the Hamiltonian can be separated in two Hamiltonians for spin up and down, respectively (see Refs. [268, 269, 273, 274, 285, 281] and Refs. therein),
@@
Figure 11: MFT result for the spatial magnetization profile, $m_{zi}(\mathbf{r}_{i})$, in a moiré cell (unit cell) of TBG at $\theta=1.30^{\circ}$ ($N=7804$) with $U=t_{0}=2.7$ eV: (a) Antiferromagnetic (AF) solution ($M_{tot}=0$, with intralayer AF order and interlayer FERRO order), and (b) almost ferromagnetic FERRO solution ($M_{tot}\neq 0$, with a strong FERRO magnetization in the AA stacking region and a small opposite magnetization everywhere else). The AA stacking regions are at the corners of the moiré cell.
@@
To perform a more systematic study of the effect of parameters on magnetization, we had also used a rescaled TB model proposed by L. A. Gonzalez-Arraga et al. [255] that simulates the magic-angle TBG flat band by a twisted bilayer containing only 869 atoms. It allows a more systematic study presented in Ref. [254]. Figures 10 and 11 show MFT results for actual TBGs, without the rescaled TB model. The main results of this work are summarized below.
@@
Consider the critical Coulomb interaction $U_{C}$ above which magnetic states emerge. In monolayer graphene, mean-field theory yields $U_{Cm}\approx 2.23t_{0}$ [286] and $U_{Cm}\approx 2.09t_{0}$ [254] for only nearest-neighbor hopping and longer-range hoppings included, respectively. For TBG, and as excepted with flat bands, i.e., for twist angles close to the magic angle, $U_{C}$ is smaller than what would be required to render an isolated layer antiferromagnetic, i.e., $U_{C}\ll U_{Cm}$ (Fig. 10(a)). These static MFT results were qualitatively confirmed by dynamical mean-field (DMFT) calculations [254] performed for the rescaled TB model. Moreover, as shown in Fig. 10(a), $U_{C}(\theta=1.30^{\circ})<U_{C}(\theta=1.08^{\circ})$, while the velocity of the corresponding flat bands are both very low (Fig. 7). This shows the importance of the moiré cell size (and then the size of the AA region) in the value of $U_{C}$. There is therefore a complex competition between the low band dispersion and low moiré pattern size for the appearance of magnetism in TBG [254].
@@
The real-space magnetic order for $U_{C}\leq U\leq U_{Cm}$ is a delicate issue. For non half filling, ferromagnetic states have been observed in experiments on doped TBG [256], and predicted theoretically [287, 288, 289, 290]. In our calculation, due to the large number of atoms of a moiré cell of magic-angle TBG ($N$ about 10000 orbitals), it is difficult to be sure that the convergence of the self-consistent process reaches the most stable magnetic state (lowest total energy). Indeed, for half filling (undoped system) and $U_{C}<U<U_{Cm}$, depending on initial conditions, the self-consistent calculation converges to an antiferromagnetic (AF) solution [254], or a ferromagnetic (FERRO) solution (Figs. 10 and  11). Note that the AF solution is in fact AF for intralayer order and FERRO for interlayer order (Fig. 11(a)). The energy difference between the AF and FERRO solutions is very small. For example, considering the TBG with $\theta=1.3^{\circ}$ (Fig. 10(b)), this difference is less than 0.1 meV per orbital, which seems not really significant to permit a definite conclusion. Let us just say that, according to this result, the FERRO solution seems more stable for small $U$, typically $U<1.4t_{0}$, and the AF solution seems to be more stable for larger $U$, $1.4t_{0}<U\lesssim U_{C}$. In both solutions, the magnetization appears in the AA region of the moiré (Fig.  11), as the flat-band states are located in the AA regions, and only flat bands are magnetic (Fig. 10(d,e,f)). As expected, the AF solution is characterized by a small gap, while the FERRO solution is metallic with a splitting of up and down bands. It is also interesting to note that, for $U<U_{Cm}$, the total magnetization of the FERRO solution is $M_{tot}=2$ electrons per moiré cell, which corresponds exactly to the magnetization of the 4 flat bands at half-filling (see dashed line Fig. 10(b)).
@@
The above calculations assume that the periodicity of the magnetic order is equivalent to the periodicity of a unit cell containing a single moiré cell, thus allowing only the studies of magnetic order within a moiré cell. Moreover, an analysis of the noninteracting susceptibility shows [254] that in real space the magnetic order inside each moiré cell should be modulated from one moiré cell to another with a wave vector $\bf q$ (${\bf q}\neq 0$).
Several theoretical studies have been performed for variable fillings (above half filling) [287, 288, 289, 290]. A delicate point is the fact that beyond half-filling it is possible for local non-collinear magnetic moments to stabilize the magnetic energy, as is the case for the honeycomb lattice for quarter doping per orbital, i.e., when $E_{F}$ is at a Van Hove singularity [291, 292, 293]. To calculate the magnetism in these cases, the Hubbard model must be treated within the framework of the Hartree-Fock approximation.
@@ 2.4 Twisted bilayer MoS2
Figure 12: (Center) TB band dispersion of twisted bilayer MoS2 with rotation angle $\theta=1.61^{\circ}$. (Surrounding panels) Average weight of the eigenstates at K of the flat bands around the gap in real space. The color scale shows the weight of the eigenstate on each $d_{z^{2}}$ orbital of the Mo atoms. The sum of these weights is more than 98% and 95% of each state for the valence and conduction band, respectively. Black lines show the unit cell containing 2522 Mo atoms. AA stacking regions are at the corners of this cell. TBGs are rigid, i.e., without atomic relaxation. From [152, 294].
@@
The TBGs that we have presented so far correspond to twisted massless 2D materials. It is also interesting to consider the appearance of flat bands in twisted massive 2D materials. That is obtained in many systems based on a honeycomb lattice, when equivalence between the two sublattice is removed, thus resulting in a gap at $E_{F}$. Among the many examples now existing of such a twisted bilayer semiconductor, the broad family of transition metal dichalcogenides (TMDs) [143, 144, 145], which offers a wide variety of possible rotationally stacked bilayer systems, has also prompted numerous experimental [295, 296, 297, 298, 299, 300, 301, 302, 151, 303, 304] and theoretical [305, 306, 307, 308, 309, 310, 311, 302, 312, 313, 314, 315, 316, 317, 318, 153, 319, 320, 321, 322, 304, 323, 324] studies to understand confined moiré pattern states in semiconductor materials. Many of these studies analyze the interlayer distances, the possible atomic relaxation, the transition from a direct band gap in the monolayer system to an indirect band gap in bilayer systems, and more generally the effect of interlayer coupling in those twisted bilayer 2D systems at various rotation angles $\theta$. For small values of $\theta$, the emergence of flat bands has been established [312] from first-principles density functional theory calculations in twisted bilayer MoS2 (tb-MoS2), and observed in a 3 twisted bilayer WSe2 sample by using scanning tunneling spectroscopy [303]. It has been also shown numerically [317] that lithium intercalation in tb-MoS2 increases interlayer coupling and thus promotes flat bands around the gap. There is also experimental and theoretical evidence that moiré patterns may give rise to confined states due to the mismatch of the lattice parameters in MoS2-WSe2 heterobilayers [151, 324]. Correlated electronic phases have been observed experimentally in twisted bilayer WSe2 over twist angles ranging from 4 to 5.1[151], and recently even superconductivity [156].
@@
Figure 13: DFT band dispersion of (red) rigid (i.e., non-relaxed), and (black) relaxed twisted bilayer MoS2 with rotation angle $\theta=5.09^{\circ}$. The zoom show the highest valence bands. Calculations are performed with the ABINIT software [325] with LDA pseudopotential. For calculation details and data see Ref. [326]. Similar results have been published previously see Refs. [312, 327].
@@
Most theoretical investigations of the electronic structure of bilayer MoS2 are density-functional theory (DFT) studies [295, 296, 297, 300, 305, 306, 307, 308, 309, 298, 310, 311, 301, 312, 328, 329, 330, 331, 320, 327, 326] with eventually a Wannier wave function analysis [306]. Those approaches provide interesting results, but they do not allow a systematic analysis of the electronic structure as a function of the rotation angle $\theta$, in particular for small angles, i.e., very large moiré cells, for which DFT calculations are not feasible. Several tight-binding (TB) models, based on Slater-Koster (SK) parameters [238], have been proposed for monolayer MoS2 [332, 333, 334, 335, 336] and multi-layer MoS2 [332, 305, 306, 334]. Following these efforts, we propose [294, 152] a SK set of parameters that match correctly the DFT bands around the gap of tb-MoS2 with rotation angles $\theta>5^{\circ}$. This SK-TB model, with the same parameters, is then used for smaller angles, in order to describe the confined moiré pattern states. We thus show that, for $\theta\lesssim 6^{\circ}$, the valence band with the highest energy is separated from the other valence states by a minigap of a few meV. In addition, the width of this band decreases as $\theta$ decreases so that the average velocity of these electronic states reaches 0 for $\theta\lesssim 2^{\circ}$ such that flat bands emerge at these angles (Fig. 12). This is reminiscent of the vanishing of the velocity at certain “magic” rotation angles in bilayer graphene (Sec. 2.3.3) except that in the case of bilayer MoS2 it arises for an interval of angles. Other minigaps and flat bands are also found in the conduction band. The confined states that are closest to the gap are localized in the AA stacking regions of the moiré pattern, as in twisted bilayer graphene. Note that there are two types of stacking for tb-MoS2 and both yield flat bands for different, but fairly comparable values of the angles [152].
@@
More recent results [312, 320, 327, 326, 324] have shown that atomic relaxation are important in tb-TMD, even for not too small twist angles. For angles $\theta$ close to 5, the atomic relaxations are mainly out-of-plane relaxations and they are well described by a simple model of atomic positions in a moiré cell [326]. As a result, the flat bands are still present, but they are modified compared to the non-relaxed system, and their degeneracy is also modified (Fig. 13). We are therefore working on a Slater-Koster TB model that will correctly simulate these relaxation effects. It should be noted that, to our knowledge, there is as yet no SK-TB model that correctly simulates the effect of this out-of-plane relaxation on the electronic structure. For smaller angles [312, 320, 327, 149], typically 2 or less, in-plane relaxation is very important and has been shown numerically to strongly modify the size of the AA and AB stacking regions. The less energy-efficient AA stacking zone is greatly reduced in favor of the AB and BA zones.

### v2

@@ 2.3.5 Heterostrain and atomic relaxation
The angle of rotation $\theta$ between the two graphene layers is the main parameter for the existence of flat moiré bands, however for angles close to the magic angle, other parameters, such as atomic relaxation and strain, are also very important and can strongly modify the band structure. There are numerous studies on atomic relaxation and the effect of strain in TBGs, which we cannot present here as they would require a review article devoted to them. We will only mention a few general aspects that we consider important in the context of this review.
Many experimental and theoretical studies have discussed the importance of atomic relaxation Uchida14 , Wijk15 , Dai16 , Nam17 , Gargiulo17 , Choi18 , Angeli18 , Lin18 , Lin18b , Lin18c , Koshino18 , Lucignano19 , Yoo19 , Guinea19 , Liu21 , Nguyen21 , Yananose21 , Kazmierczak21 , Leconte22 , Ezzi24 , Ceferino24 . Indeed, the presence of different stacking zones AA, AB, SP, etc. in a moiré pattern, which do not have the same stability in relation to each other Gargiulo17 , strongly suggests that atomic reconstruction (or atomic relaxation) relative to rigidly TBG is likely, and that it will be all the stronger as the size of the moiré pattern (Eq. (18)) increases. It is useful to distinguish between atomic relaxation in the graphene layers (in-plane relaxation) and perpendicular to the graphene layers (out-of-plan relaxation or $z$-modulation). The latter has been well described Gargiulo17 , Koshino18 and mainly stems from the fact that the interlayer distance is not the same for stable AA and AB stacking structures. It can modify the electronic structure and it is often argued that it is the cause of minigaps that isolate the flat bands. However, it should be noted that such minigaps can exist without out-of-plane relaxation (see, e.g., Fig. 7(a)). Furthermore, to our knowledge, the effect of this relaxation has not been well studied, in a systematic way, by DFT calculations, which are difficult to perform for such large systems. Finally, it should be noted that SK-TB models have not really been validated to account for this out-of-plane relaxation (Sec. 2.3.2). However, the SK-TB model correctly accounts for the effects of the in-plane relaxation. In the case of graphene twisted bilayers, it has been shown that this effect is very weak for angles close to a first magic angle ($\theta\simeq 1^{\circ}$) and therefore has no significant effect on the electronic structure. However, for angles smaller than the first magic angle, several studies show the importance of in-plane reconstruction Gargiulo17 , Guinea19 , Nguyen21 with a reduction in AA zones in favor of AB zones, which are energetically more stable. For a very large moiré of several millions of atoms, this atomic reconstruction is stronger and can lead to new phenomena such as giant swirls associated with new types of electronic localization Mesple23 .
@@
However, in the case of TBG on SiC close to the first magic angle, a detailed study by Huder et al. Huder18 and Mesple et al. Mesple21 shows that the effect of “heterostrain,” i.e., an in-plane uniaxial deformation of one layer with respect to the other, is more important in many samples than the atomic relaxation. It is now well established that the heterostrain is present in many samples of TBG, although not well controlled experimentally, and it has strong consequences for the electronic properties of flat bands Huder18 , Bi19 , Mesple21 , Kerelsky2019 , Gao21 , Pantaleon21 , Mannai21 , Parker21 , XIONG2023129048 , Escudero24 , Yu2025 . To study it in a quantitative way, it is then possible to calculate the local electronic structure from our SK-TB model (Sec. 2.3.2), without other adjustable parameters. This approach has allowed to study several samples that have have been studied from STM images in the literature (see Refs. Huder18 , Mesple21 , Huder17_these , FlorieThese and Refs. therein). The main result is that for TBGs with angles close to the first magic angle, a small amount of heterostrain is an ubiquitous factor in strongly modifying the electronic structure of the low-energy flat bands. For example, the uniaxial heterostrain can induce the appearance of a third peak in the low-energy DOS in addition to two van Hove singularities Huder18 . Moreover heterostrain modifies also the electron-electron interactions in many samples that have Fermi levels in the flat bands Mesple21 . It therefore appears that for the first magic-angle TBGs studied experimentally, the heterostrain is a determining parameter, more important than the atomic relaxations.
@@ 2.3.6 Magnetism
Figure 10: MFT results for TBG: (a) Magnetization ($\mu_{B}$) at zero temperature ($T=0$) as a function of $U$ at rotation angles $\theta=1.085^{\circ}$ (bilayer (30,31), $N=11164$), $\theta=1.297^{\circ}$ ((25,26), $N=7804$), and $\theta=1.47^{\circ}$ ((22,23), $N=6076$), respectively. NN is the number of orbitals per moiré cell. The monolayer case is calculated with similar conditions (similar supercell). For a monolayer, the critial $U$ above which the solution is magnetic is $U_{Cm}\approx 2.09t_{0}$, where $t_{0}=2.7$ eV is the first intralayer coupling. The other panels are for $\theta=1.297^{\circ}$ (25,26) TBG: (b) Magnetization per moiré cell: (solid line) $M_{tot}^{*}(U)$ for the AF and FERRO solutions, and (dashed line) $M_{tot}(U)$ for FERRO ($M_{tot}=0$ for AF), see Eq. (31). (c) $M^{*}_{tot}(U=t_{0}=2.7\,{\rm eV})$ versus temperature $T$ for the AF solution. (d) Up and down local DOSs on an orbital with up magnetization at the center of the AA stacking region for $U=t_{0}=2.7$ eV for the AF solution. (e) and (f) Bands around $E_{F}$ of the AF and FERRO solutions for $U=t_{0}=2.7$ eV. Up and down bands are degenerate for the AF solution. SK-TB model is the one which results in an intraband monolayer velocity $V_{MLG}=0.79\cdot 10^{6}$ ms-1 (Sec. 2.3.2). TBGs are rigid, i.e., without atomic relaxation. For details of the MFT calculations see Ref. Vahedi21_tBLG .
The large reduction in kinetic energy in flat bands increases the relative importance of interactions, making bilayer systems much more susceptible to correlation effects than a single layer Luis17 , Cao18a , Sharpe19 , Klebl19 , Goodwin19 , klebl2020 , Cea20 , Chen22 , JimenoPozo23 , Dale23 , Wilhelm23 , Rai24 , Jiang25_gap_miniband , Kwan25 (see also Vahedi21_tBLG and Refs. therein). One example is the magnetization of edge states in graphene nanoribbons JPSJ.65.1920 , JPSJ.67.2089 , PhysRevB.59.8271 , PhysRevLett.101.246803 , Feldner10 , Yazyev2010 , Wakabayashi10 , Tao11 , Feldner11 , FeldnerThese , Wakabayashi13 , TMD nanoribbons vahedi21_TMDedge , or graphene nanoflakes PhysRevLett.99.177204 , Bhowmick2008 , Viana-Gomes2009 , Feldner10 , Yazyev2010 , FeldnerThese , Roy2014 , Valli2016 , Thu20 , Thu_these , Phung22 . However, in TBGs the flat bands at the origin of strong electronic correlations are not due to defects or edges, but to the particular symmetry of the moiré pattern. We have first investigated Vahedi21_tBLG the question of magnetic instabilities at half filling of TBG using our tight-binding description of the non-interacting bilayer systems (Eq. (15)) to which we add an on-site Hubbard interaction $U$ in order to model the Coulomb repulsion between electrons of opposite spin Hubbard1963 ,
@@
where $\sigma=\,\uparrow$ ($\downarrow$) is the spin up (down), $\hat{H}_{0\sigma}$ is the non-interacting SK-TB Hamiltonian for each spin (Eq. (19)), and $n_{i\sigma}=c_{i\sigma}^{{\dagger}}c_{i\sigma}$ is the number operator at site $i$ and $\mu_{c}$ the chemical potential. In the static mean-field (MFT) approximation, assuming that local magnetic moments are collinear to the $z$-direction (Hartree approximation), the Hamiltonian can be separated in two Hamiltonians for spin up and down, respectively (see Refs. Feldner10 , Yazyev2010 , FeldnerThese , Wakabayashi13 , Marcin19 , Thu20 and Refs. therein),
@@
Figure 11: MFT result for the spatial magnetization profile, $m_{zi}(\mathbf{r}_{i})$, in a moiré cell (unit cell) of TBG at $\theta=1.297^{\circ}$ ((25,26) $N=7804$) with $U=t_{0}=2.7$ eV: (a) Antiferromagnetic (AF) solution ($M_{tot}=0$, with intralayer AF order and interlayer FERRO order), and (b) almost ferromagnetic FERRO solution ($M_{tot}\neq 0$, with a strong FERRO magnetization in the AA stacking region and a small opposite magnetization everywhere else). The AA stacking regions are at the corners of the moiré cell.
@@
To perform a more systematic study of the effect of parameters on magnetization, we had also used a rescaled TB model proposed by L. A. Gonzalez-Arraga et al. Luis17 that simulates the magic-angle TBG flat band by a twisted bilayer containing only 869 atoms. It allows a more systematic study presented in Ref. Vahedi21_tBLG . Figures 10 and 11 show MFT results for actual TBGs, without the rescaled TB model. The main results of this work are summarized below.
@@
Consider the critical Coulomb interaction $U_{C}$ above which magnetic states emerge. In monolayer graphene, mean-field theory yields $U_{Cm}\approx 2.23t_{0}$ Sorella1992 and $U_{Cm}\approx 2.09t_{0}$ Vahedi21_tBLG for only nearest-neighbor hopping and longer-range hoppings included, respectively. For TBG, and as excepted with flat bands, i.e., for twist angles close to the magic angle, $U_{C}$ is smaller than what would be required to render an isolated layer antiferromagnetic, i.e., $U_{C}\ll U_{Cm}$ (Fig. 10(a)). These static MFT results were qualitatively confirmed by dynamical mean-field (DMFT) calculations Vahedi21_tBLG performed for the rescaled TB model. Moreover, as shown in Fig. 10(a), $U_{C}(\theta=1.30^{\circ})<U_{C}(\theta=1.08^{\circ})$, while the velocity of the corresponding flat bands are both very low (Fig. 7). This shows the importance of the moiré cell size (and then the size of the AA region) in the value of $U_{C}$. There is therefore a complex competition between the low band dispersion and the moiré pattern size for the appearance of magnetism in TBG Vahedi21_tBLG .
@@
The real-space magnetic order for $U_{C}\leq U\leq U_{Cm}$ is a delicate issue. Away from half filling, ferromagnetic states have been observed in experiments on doped TBG Sharpe19 , and predicted theoretically Wu19_FERRO , Seo19 , pons2020flatband , Vidarte24 . In our calculation, due to the large number of atoms of a moiré cell of magic-angle TBG ($N$ about 10000 orbitals), it is difficult to be sure that the convergence of the self-consistent process reaches the most stable magnetic state (lowest total energy). Indeed, for half filling (undoped system) and $U_{C}<U<U_{Cm}$, depending on initial conditions, the self-consistent calculation converges to an antiferromagnetic (AF) solution Vahedi21_tBLG , or a ferromagnetic (FERRO) solution (Figs. 10 and  11). Note that, for $U=t_{0}$, the AF solution is in fact AF for intralayer order and FERRO for interlayer order (Fig. 11(a)). The energy difference between the AF and FERRO solutions is very small. For example, considering the TBG with $\theta=1.3^{\circ}$ (Fig. 10(b)), this difference is less than 0.1 meV per orbital, which seems not really significant to permit a definite conclusion. Let us just say that, according to this result, the FERRO solution seems more stable for small $U$, typically $U<1.4t_{0}$, and the AF solution seems to be more stable for larger $U$, $1.4t_{0}<U\lesssim U_{Cm}$. In both solutions, the magnetization appears in the AA region of the moiré (Fig.  11), as the flat-band states are located in the AA regions, and only flat bands are magnetic (Fig. 10(d,e,f)). As expected, the AF solution is characterized by a small gap, while the FERRO solution is metallic with a splitting of up and down bands. It is also interesting to note that, for $U<U_{Cm}$, the total magnetization of the FERRO solution is $M_{tot}=2$ electrons per moiré cell, which corresponds exactly to the magnetization of the 4 flat bands at half-filling (see dashed line Fig. 10(b)).
@@
The above calculations assume that the periodicity of the magnetic order is equivalent to the periodicity of a unit cell containing a single moiré cell, thus allowing only the studies of magnetic order within a moiré cell. Moreover, an analysis of the noninteracting susceptibility suggests Vahedi21_tBLG that in real space the magnetic order inside each moiré cell could be modulated from one moiré cell to another with a wave vector $\bf q$ (${\bf q}\neq 0$). Further investigations are needed to clarify this point.
Several theoretical studies have been performed for variable fillings (above half filling) Wu19_FERRO , Seo19 , pons2020flatband , Vidarte24 . A delicate point is the fact that beyond half filling it is possible for local non-collinear magnetic moments to stabilize the magnetic energy, as is the case for the honeycomb lattice for quarter doping per orbital, i.e., when $E_{F}$ is at a Van Hove singularity Wang12_MC_graphene1-4 , Jiang14 , Maxime25 . To calculate the magnetism in these cases, the Hubbard model must be treated within the framework of the Hartree-Fock approximation.
@@ 2.4 Twisted bilayer MoS2
Figure 12: (Center) TB band dispersion around the gap of rigid twisted bilayer MoS2 with rotation angle $\theta=1.61^{\circ}$. (Surrounding panels) Average weight of the eigenstates at K of the flat bands around the gap in real space. The color scale shows the weight of the eigenstate on each $d_{z^{2}}$ orbital of the Mo atoms. The sum of these weights is more than 98% and 95% of each state for the valence and conduction band, respectively. Black lines show the unit cell containing 2522 Mo atoms. AA stacking regions are at the corners of this cell. TBGs are rigid, i.e., without atomic relaxation. From Venky20 , Venky_these .
@@
The TBGs that we have presented so far correspond to twisted massless 2D materials. It is also interesting to consider the appearance of flat bands in twisted massive 2D materials. That is obtained in many systems based on a honeycomb lattice, when equivalence between the two sublattices is removed, thus resulting in a gap at $E_{F}$. Among the many examples now existing of such a twisted bilayer semiconductor, the broad family of transition metal dichalcogenides (TMDs) Wang15_reviewTMD , Liu15_reviewTMD , Duong17_review , which offers a wide variety of possible rotationally stacked bilayer systems, has also prompted numerous experimental vanderZande14 , Liu14 , Huang14 , Huang16 , Yeh16 , Zhang17 , Trainer17 , Wilson17 , Pan18 , Zhang20_tTMD , Graham2021 and theoretical Roldan14b , Fang15 , Cao15 , Wang15 , Constantinescu15 , Tan16 , Lu17 , Wilson17 , Naik18 , Conte19 , Maity19 , Tang20 , Wu20 , Lu20 , Pan20 , Zhan20 , Zhang20 , Naik20 , Xian21 , Angelie21 , Graham2021 , PhysRevB.110.235410 , Soltero24 studies to understand confined moiré pattern states in semiconductor materials. Many of these studies analyze the interlayer distances, the possible atomic relaxation, the transition from a direct band gap in the monolayer system to an indirect band gap in bilayer systems, and more generally the effect of interlayer coupling in those twisted bilayer 2D systems at various rotation angles $\theta$. For small values of $\theta$, the emergence of flat bands has been established Naik18 from first-principles density functional theory calculations in twisted bilayer MoS2 (tb-MoS2), and observed in a 3 twisted bilayer WSe2 sample by using scanning tunneling spectroscopy Zhang20_tTMD . It has been also shown numerically Lu20 that lithium intercalation in tb-MoS2 increases interlayer coupling and thus promotes flat bands around the gap. There is also experimental and theoretical evidence that moiré patterns may give rise to confined states due to the mismatch of the lattice parameters in MoS2-WSe2 heterobilayers Pan18 , Soltero24 . Correlated electronic phases have been observed experimentally in twisted bilayer WSe2 over twist angles ranging from 4 to 5.1Pan18 , and recently even superconductivity Guo25 .
@@
Figure 13: DFT band dispersion of (red) rigid (i.e., non-relaxed), and (black) relaxed twisted bilayer MoS2 with rotation angle $\theta=5.09^{\circ}$. The zoom show the highest valence bands. Calculations are performed with the ABINIT software Gonze16 with LDA pseudopotential. For calculation details and data see Ref. venky23 . Similar results have been published previously see Refs. Naik18 , Vitale21 .
@@
Most theoretical investigations of the electronic structure of bilayer MoS2 are density-functional theory (DFT) studies vanderZande14 , Liu14 , Huang14 , Zhang17 , Roldan14b , Fang15 , Cao15 , Wang15 , Constantinescu15 , Huang16 , Tan16 , Lu17 , Trainer17 , Naik18 , Debbichi14 , Tao14 , He14 , sun2020effects , Naik20 , Vitale21 , venky23 with eventually a Wannier wave function analysis Fang15 . Those approaches provide interesting results, but they do not allow a systematic analysis of the electronic structure as a function of the rotation angle $\theta$, in particular for small angles, i.e., very large moiré cells, for which DFT calculations are not feasible. Several tight-binding (TB) models, based on Slater-Koster (SK) parameters Slater54 , have been proposed for monolayer MoS2 Cappelluti13 , Rostami13 , Zahid13 , Ridolfi15 , SilvaGuillen16 and multi-layer MoS2 Cappelluti13 , Roldan14b , Fang15 , Zahid13 . Following these efforts, we propose Venky_these , Venky20 a SK set of parameters that match correctly the DFT bands around the gap of tb-MoS2 with rotation angles $\theta>5^{\circ}$. This Hamiltonian takes into account the five $d$ orbitals of the Mo atoms and the three $p$ orbitals of the S atoms. This SK-TB model, with the same parameters, is then used for smaller angles, in order to describe the confined moiré pattern states. We thus show that, for $\theta\lesssim 6^{\circ}$, the valence band with the highest energy is separated from the other valence states by a minigap of a few meV. In addition, the width of this band decreases as $\theta$ decreases so that the average velocity of these electronic states reaches 0 for $\theta\lesssim 2^{\circ}$ such that flat bands emerge at these angles (Fig. 12). This is reminiscent of the vanishing of the velocity at certain “magic” rotation angles in bilayer graphene (Sec. 2.3.3) except that in the case of bilayer MoS2 it arises for an interval of angles. Other minigaps and flat bands are also found in the conduction band. The confined states that are closest to the gap are localized in the AA stacking regions of the moiré pattern, as in twisted bilayer graphene. Note that there are two types of stacking for tb-MoS2 and both yield flat bands for different, but fairly comparable values of the angles Venky20 .
@@
More recent results Naik18 , Naik20 , Vitale21 , venky23 , Soltero24 have shown that atomic relaxation is important in tb-TMD, even for not too small twist angles. For angles $\theta$ close to 5, the atomic relaxations are mainly out-of-plane relaxations and they are well described by a simple model of atomic positions in a moiré cell venky23 . As a result, the flat bands are still present, but they are modified compared to the non-relaxed system, and their degeneracy is also modified (Fig. 13). We are therefore working on a Slater-Koster TB model that will correctly simulate these relaxation effects. It should be noted that, to our knowledge, there is as yet no SK-TB model that correctly simulates the effect of this out-of-plane relaxation on the electronic structure. For smaller angles Naik18 , Naik20 , Vitale21 , Walet21 , typically 2 or less, in-plane relaxation is very important and has been shown numerically to strongly modify the size of the AA and AB stacking regions. The less energy-efficient AA stacking zone is greatly reduced in favor of the AB and BA zones.